A family has 8 girls and 4 boys. A total of 3 children must be chose to speak on the behalf of the family at a local benefit. What is the probability that no girls and 3 boys will be chosen?
Answer by stanbon(75887)

A family has 8 girls and 4 boys. A total of 3 children must be chose to speak on the behalf of the family at a local benefit. What is the probability that no girls and 3 boys will be chosen?
----
# of ways to succeed: 4C3 = 4
# of random triples: 8C3 = (8*7*6)/(1*2*3) = 56
---
P(3 boys chosen) = 4/56 = 1/14
----------------------------------
Cheers,
Stan H.
===================